### v1

@@ 4.2 SIMULATION TOOLS AND SETTINGS
Figure 11: The number of neighbors vs. the number of vehicles/km for (a) Highway Scenarios and (b) Urban Scenarios.
These settings allow evaluation of the trade-offs between communication overhead introduced by the proposed protocol and the reliability, timeliness, and channel efficiency of vehicular communication in both highway and urban environments. We ran our simulations on High Performance Computing (HPC) resources supported by the University of Arizona. We dedicated one node with one core for each scenario (Different simulation scenarios explained in 4.3). Standard partition on Puma cluster was used which uses CPU only. We used 6 GBytes memory per CPU. Matlab r2023b was used to run the WiLabV2XSim simulator.
4.3 SIMULATION SCENARIOS
Each simulation lasted one minute and was run for both highway and urban scenarios. In the highway case, vehicles traveled along a 5000 m bidirectional road with four lanes per direction, an average speed of 120 km/h, and a standard deviation of 20 km/h. In the urban case, traffic was modeled on a $2\times 2$ grid comprising four blocks, each 250 m long and wide, with four lanes in both vertical and horizontal directions. For both the highway and the urban scenarios, the lane width is 4 meters.The mean speed was 60 km/h with a standard deviation of 15 km/h. Simulations were run using DSRC and LTE-V2X communication protocols sending either the minimum sized packet or the average sized packet. We ran the simulations with 3 different vehicle densities in both the highway and the urban scenarios. For the highway scenario, we chose 100, 200, 300 vehicles/km densities while for the urban scenario, we chose 300, 400, 500 vehicles/km densities. Simulations for each combination of environment, communication protocol, packet size and vehicle density were repeated 10 times with different starting seed and the outputs were averaged.
@@ 4.4 SIMULATION RESULTS
This section presents and analyzes the results of our simulation-based performance evaluation. The objective is to quantify the communication overhead introduced by our Zero Trust D-IM protocol and assess its impact on critical vehicular network metrics such as Packet Reception Rate (PRR), Channel Busy Ratio (CBR), and latency. The following figures present the key findings across various scenarios, demonstrating that our framework provides strong security guarantees with minimal performance trade-offs suitable for real-world deployment.
From Figure 7, we can conclude that increasing the number of vehicles/km results in decreasing PRR. Also, for a given vehicle density, PRR decreases with increasing the distance. Moreover, we can clearly find that PRR is higher in highway scenario than it is for urban scenario. This can be explained by the fact that in urban scenarios there are many obstacles, i.e. buildings that hinder reception of packets.
Our implementation of the zero trust blockchain based identification protocol poses an overhead in terms of the size of packets exchanged. Comparing PRR when the average size required by our implementation is used for the simulation packet size to the case in which the simulation packet size is set to the minimum packet size, we can see in 8 that this overhead results in a reduction of around 7.5% in PRR for urban scenarios and reduction of 29% in PRR for highway scenario with 300 vehicles/km using LTE-V2X RAT. For the cases where DSRC RAT is used, a reduction of around 16% in PRR for urban scenarios and reduction of 72% in PRR. The huge reduction in PRR when DSRC RAT is caused by the use of CSMA/CA which in case of larger packets being transmitted, results in higher back-off rates.
@@
We, also, observe that PRR is lower when DSRC is used than when LTE-V2X is used. This is because DSRC uses carrier sense multiple access/ collision avoidance (CSMA/CA) for media access which implements "listen-before-talk" strategy, while LTE-V2X mode 4, that enables peer-to-peer communication adopted in our simulations, uses Sensing-Based Semi-Persistent Scheduling (SB-SPS) for media access. SB-SPS also uses a "listen-before-talk" strategy and a periodic reservation system to manage channel access and reduce collisions. In SB-SPS, once a resource is selected, the vehicle reserves it for a specified number of consecutive transmissions at a regular interval. This semi-persistent reservation is more efficient than a full random access protocol for every packet resulting in higher PRR.
Figure 9 shows the relationship between the CBR and the number of vehicles per km for both highway and urban scenarios. We can observe that increasing the number of vehicles/km increases the CBR for both environments. We also observe that the CBR is higher in the urban scenarios than in the highway scenarios. This is due to that fact that urban environments tend to have more vehicles than highway environments, resulting in more traffic and higher CBR. Moreover, we can observe that the overhead of our implemented protocol leads to an increase of less than 11% in CBR when LTE-V2X is used.
@@
Observing Figure 10, we can find that increasing the number of vehicles/km slightly increases data age for all scenarios. We also observe that data age is slightly higher in urban scenarios than in highway scenarios. This is due to that fact that urban environments are more congested with packets. Moreover, we can observe that the overhead of our implemented protocol leads to an increase of less than 0.06 seconds in data age for LTE-V2X simulations and less than 0.3 sec in DSRC simulations.
@@
In summary, our simulation results validate the practicality and efficiency of the Zero Trust D-IM protocol. The framework’s overhead on network performance is minimal and well within acceptable limits for safety-critical V2X applications. Our findings show that even in dense traffic, the solution maintains robust network metrics.
@@ 5 CONCLUSIONS AND FUTURE WORK
In this paper, we introduced a novel Zero Trust–based Decentralized Identity Management (D-IM) framework for autonomous vehicles, addressing the critical security and trust challenges inherent in V2X communication. By integrating the core principles of Zero Trust Architecture with the immutable, decentralized nature of blockchain, our solution eliminates the single points of failure and trust assumptions found in traditional centralized systems like Public Key Infrastructure (PKI). We designed and implemented a blockchain-based identification and authorization protocol that ensures continuous verification of every entity, leveraging Hyperledger Iroha to provide a lightweight and efficient distributed trust anchor.
Our comprehensive experimental evaluation, conducted across both urban and highway scenarios, validates the practicality and performance of the D-IM framework. The results demonstrate that while the protocol introduces a minimal overhead, a slight reduction in Packet Reception Rate (PRR), a small increase in Channel Busy Ratio (CBR), and negligible added latency, these trade-offs are well within acceptable limits for real-time, safety critical applications. Specifically, our findings confirm that the D-IM protocol can deliver robust security guarantees, including strong resistance to impersonation and replay attacks, without compromising the efficiency or reliability of the underlying communication network. By providing a resilient, scalable identity management solution, our work provides a solid foundation for securing the next generation of connected and autonomous vehicles.

### v2

@@ 6.2 Evaluation Parameters and Scenarios
  • Update Delay or Inter-Packet Gap (IPG): Time between two consecutive successfully received beacons from the same node within the awareness range.
@@
Figure 6: (a) Highway Packet Reception Rate and (b) Urban Packet Reception Rate vs. the distance using DSRC RAT and average packet size required by our implementation (c) Highway Packet Reception Rate and (d) Urban Packet Reception Rate vs. the distance using LTE-V2X RAT and average packet size required by our implementation.
@@ 6.3 Simulation Results
This section presents and analyzes the results of our simulation-based performance evaluation and presents a scalability analysis increasing the vehicle density for each scenario. The objective is to quantify the communication overhead introduced by our Zero Trust D-IM protocol, and assess its impact on critical vehicular network metrics such as PRR, CBR, and latency, and its scalability potential. The following figures present the key findings across various scenarios, demonstrating that our framework provides strong security guarantees with minimal performance trade-offs suitable for real-world deployment.
Figure 7: (a) Highway Packet Reception Rate and (b) Urban Packet Reception Rate vs. the Number of Vehicles per km at 50m.
@@
Figure 6 shows the PRR vs distance with different communication protocols, DSRC AND C-V2X, for both urban and highway environments. The curves show that increasing the number of vehicles per km results in decreasing the PRR. Also, for a given number of vehicles per km, the PRR decreases with increasing the distance. Moreover, from the figure, we clearly find that PRR is higher in highway scenario than it is for urban scenario. This can be explained by the fact that in urban scenarios there are many obstacles, i.e. buildings that hinder the reception of packets. We, also, observe that the PRR is lower when DSRC is used than when LTE-V2X is used. This is because DSRC uses carrier sense multiple access/ collision avoidance (CSMA/CA) for media access which implements "listen-before-talk" strategy, while LTE-V2X mode 4, that enables peer-to-peer communication adopted in our simulations, uses Sensing-Based Semi-Persistent Scheduling (SB-SPS) for media access. SB-SPS also uses a "listen-before-talk" strategy and a periodic reservation system to manage channel access and reduce collisions. In SB-SPS, once a resource is selected, the vehicle reserves it for a specified number of consecutive transmissions at a regular interval. This semi-persistent reservation is more efficient than a full random access protocol for every packet resulting in higher PRR.
Our implementation of the zero trust blockchain based identification protocol poses an overhead in terms of the size of packets exchanged. To compare PRR when the average size required by our implementation is used for the simulation packet size to the case in which the simulation packet size is set to the minimum packet size, we refer to Figure 7 which presents the PRR vs the vehicle density per km with different communication protocols for both urban and highway environments. The graphs show that this overhead results in an average reduction of around 18% in PRR for urban scenarios and reduction of 6% in PRR for highway scenario using LTE-V2X RAT. For vehicle density of 200 vehicles/km, the reduction observed was of around 19% in PRR for urban scenarios and around 3% in PRR for highway scenario using LTE-V2X RAT. For the cases where DSRC RAT is used, a average reduction of around 27% in PRR for urban scenarios and average reduction of 4% in PRR for highway scenarios. For vehicle density of 200 vehicles/km, the reduction observed was of around 28% in PRR for urban scenarios and around 6% in PRR for highway scenario using DSRC RAT.
Figure 8: (a) The Highway Data Age and (b) The Urban Data Age vs. the number of vehicles per km (c) The Highway Update Delay and (d) The Urban Update Delay vs. the number of vehicles per km
@@
From Figures 8a, 8b, we can find that increasing the number of vehicles/km increases the data age for both scenarios. The graphs show that the data age is higher in the urban scenarios than in the highway scenarios. This is due to that fact that urban environments are more congested with packets. When using DSRC RAT in our D-IM, the data age is much higher because of DSRC RAT use of CSMA/CA for madia access which cause longer latencies. Moreover, we can observe that the overhead of our implemented protocol leads to an increase of less than 0.032 seconds in data age for LTE-V2X simulations and less than 0.5 sec in DSRC simulations. The same can be concluded for the inter-packet gap (update delay) in Figures 8c, 8d.
Figure 9: (a) Highway Channel Busy ratio and (b) Urban Channel Busy ratio vs. the number of vehicles per km
Figure 9 shows the relationship between the CBR and the number of vehicles per km for both highway and urban scenarios. We can observe that increasing the number of vehicles/km increases the CBR for both environments. We also observe that the CBR is higher in the urban scenarios than in the highway scenarios when DSRC RAT is used. This is due to that fact that urban environments tend to have more vehicles than highway environments, resulting in more traffic and higher CBR given the fact that DSRC RAT uses "listen before talk" strategy . The overhead of our implemented protocol leads to an average increase of less than 16% in CBR when LTE-V2X is used, and less than 27% when DSRC is used.
Figure 10: The number of neighbors vs. the number of vehicles per km for (a) Highway Scenarios and (b) Urban Scenarios.
@@
In summary, our simulation results validate the practicality and efficiency of the Zero Trust D-IM protocol. The framework’s overhead on network performance is limited and well within acceptable limits for safety-critical V2X applications. Our findings show that even in dense traffic, the solution maintains robust network metrics. Specifically, in urban scenarios, the overhead from our protocol led to a manageable PRR degradation of around 18% with LTE-V2X and 27% with DSRC. For highway scenarios, the PRR reduction was a practical 6% for LTE-V2X and 4% for DSRC. Furthermore, our implementation added a minimal increase in Channel Busy Ratio (CBR) of less than 16% for LTE-V2X (12% for urban scenarios). Most critically for real-time applications, the latency overhead was negligible, with an increase of less than 0.032 seconds for LTE-V2X and under 0.5 seconds for DSRC. The maximum overhead observed and reported in this paper is for 500 vehicle/km density which is much more than the real-world expected average vehicle density. The overhead for lower vehicle densities is much lower. These results collectively demonstrate that our blockchain-based solution can deliver robust, continuous identity verification and security guarantees without compromising the efficiency or reliability of the underlying communication network, without relying on implicit trust.
@@ 7 CONCLUSIONS AND FUTURE WORK
This paper presented D-IM, a Zero Trust-based, blockchain-enabled decentralized identity management and authentication framework for autonomous vehicle (AV) communication systems. Motivated by the limitations of perimeter-based security models and centralized certificate authorities, D-IM was designed to address a comprehensive attacker model that captures identity-based, data integrity, collusion, availability, and trust-related threats in highly dynamic V2X environments.
We first defined a set of explicit design goals aligned with Zero Trust principles, including continuous verification, mutual authentication, decentralization, privacy preservation, non-repudiation, and traceability. Building on these goals, we introduced the D-IM system architecture and protocol, which leverage a permissioned blockchain as a decentralized trust anchor for identity registration, and verification. By binding vehicle identities to cryptographic credentials stored immutably on the blockchain and enforcing identity verification prior to any data exchange, D-IM eliminates implicit trust assumptions and reduces reliance on centralized authorities.
To evaluate the security properties of the proposed framework, we conducted both qualitative and formal security analyses, including a BAN logic-based verification of the identification and authorization protocol. The analysis demonstrates that D-IM achieves mutual authentication, fresh session key establishment, and resistance to impersonation, replay, man-in-the-middle, repudiation, and single-point-of-failure attacks under the defined threat model. Attacks that cannot be fully prevented at the identity layer, such as data falsification or collusion, are effectively detected, attributed, and contained through decentralized verification and accountability mechanisms.
We further validated the practicality of D-IM through simulation-based performance evaluations in realistic urban and highway scenarios using both DSRC and C-V2X technologies. The results show that continuous identity verification and blockchain-assisted authentication introduce limited communication and latency overhead, even under dense traffic conditions. Specifically, the framework maintains acceptable packet reception ratios, channel utilization, and end-to-end latency, demonstrating that strong security and accountability guarantees can be achieved without compromising the performance requirements of real-world AV deployments.
Overall, D-IM establishes a robust security foundation for autonomous vehicle communication networks by tightly integrating Zero Trust principles with decentralized identity management. The framework successfully addresses the defined attacker model while remaining scalable and efficient, making it well suited for next-generation intelligent transportation systems.
@@
Future work will further strengthen the proposed framework through the integration of smart contract-based enforcement mechanisms and anomaly-based machine learning techniques. These extensions aim to automate misbehavior penalties, enhance detection of subtle or coordinated attacks, and improve system adaptability in evolving threat environments. Additionally, future research will explore large-scale deployments, more diverse adversarial behaviors, and deeper integration with cooperative perception and decision-making systems.